OpenStudy (anonymous):

Solve this system by the addition method. 5x+y=26 x+5y=-14 Put answer in ordered pair form.

OpenStudy (anonymous):

well what we can do is look for the slope so we need to subtract 5x and leave the y alone so y=-5x+26 so the slope is -5x

OpenStudy (anonymous):

then we look for the y intercept so is like this 

OpenStudy (anonymous):

so now we know the first ordered pair (5.2,26)

OpenStudy (anonymous):

now we do the same with the other equation

OpenStudy (anonymous):

x+5y=-14 subtract x 5y=x-14 now we divide by 5 so we can find y y=x/5-2.8

OpenStudy (anonymous):